in a certain class, test scores average out to 60, with a standard deviation of 20, as do scores on the final. The correlation between test scores and the final scores is about 0.5. Estimate the average final score for the students whose test score was: 60

A. 60 B. 70 C. 50 D. 80
Answer #1

Estimate the average final score for the students whose test score was: 60

60 is the student test score + 1 standard deviation

60 = 60+20 = 60 + 1 SD

now multiple the correlation coefficient (0.5) by 1

0.50 * 1 = 0.50

then multiple the 0.50 by the SD of 20

0.50 * SD = 0.50 * 20 = 10

now add 10 to 60

60 + 10 = 70

70 is the final score (Option B)